# Overcoming Slow Decision Frequencies in Continuous Control: Model-Based Sequence Reinforcement Learning for Model-Free Control

### v3

@@ Abstract
Reinforcement learning (RL) is rapidly reaching and surpassing human-level control capabilities. However, state-of-the-art RL algorithms often require timesteps and reaction times significantly faster than human capabilities, which is impractical in real-world settings and typically necessitates specialized hardware. We introduce Sequence Reinforcement Learning (SRL), an RL algorithm designed to produce a sequence of actions for a given input state, enabling effective control at lower decision frequencies. SRL addresses the challenges of learning action sequences by employing both a model and an actor-critic architecture operating at different temporal scales. We propose a ”temporal recall” mechanism, where the critic uses the model to estimate intermediate states between primitive actions, providing a learning signal for each individual action within the sequence. Once training is complete, the actor can generate action sequences independently of the model, achieving model-free control at a slower frequency. We evaluate SRL on a suite of continuous control tasks, demonstrating that it achieves performance comparable to state-of-the-art algorithms while significantly reducing actor sample complexity. To better assess performance across varying decision frequencies, we introduce the Frequency-Averaged Score (FAS) metric. Our results show that SRL significantly outperforms traditional RL algorithms in terms of FAS, making it particularly suitable for applications requiring variable decision frequencies. Furthermore, we compare SRL with model-based online planning, showing that SRL achieves comparable FAS while leveraging the same model during training that online planners use for planning. Lastly, we highlight the biological relevance of SRL, showing that it replicates the ”action chunking” behavior observed in the basal ganglia, offering insights into brain-inspired control mechanisms.
@@ 1 Introduction
Biological and artificial agents must learn behaviors that maximize rewards to thrive in complex environments. Reinforcement learning (RL), a class of algorithms inspired by animal behavior, facilitates this learning process (Sutton and Barto, 2018). The connection between neuroscience and RL is profound. The Temporal Difference (TD) error, a key concept in RL, effectively models the firing patterns of dopamine neurons in the midbrain (Schultz et al., 1997; Schultz, 2015; Cohen et al., 2012). Additionally, a longstanding goal of RL algorithms is to match and surpass human performance in control tasks (OpenAI2019; Schrittwieser et al., 2020; Kaufmann et al., 2023a; Wurman et al., 2022a; Vinyals et al., 2019; Mnih et al., 2015)
@@
In contrast, biological agents achieve precise and seemingly fast control using much slower hardware (biological neurons). This is possible because biological agents effectively decouple the computation frequency from the actuation frequency, allowing them to achieve high actuation frequencies even with slow computational speeds. Consequently, biological agents demonstrate robust, adaptive, and efficient control.
@@ 4 Sequence Reinforcement Learning
We introduce a novel reinforcement learning model capable of learning sequences of actions (macro-actions) by replaying memories at a finer temporal resolution than the action generation, utilizing a model of the environment during training. We provide the neural basis for our algorithm in Section 6.
@@
5 Experimental Setup
@@
Figure 2: Performance of SRL and SAC at different Action Sequence Lengths (ASL) for the Hopper environment. All markers are averaged over 5 trials, with the error bars representing standard error. The Frequency-Averaged Score (FAS) is calculated as the area under the curve of this plot.
Transitioning from simulation to real-world implementation (Sim2Real) in control systems is challenging because deployment introduces computational stochasticity, leading to variable sensor sampling rates (throughput) and inconsistent end-to-end delays from sensing to actuation (Sandha et al., 2021). This gap is not captured by the mean reward or return that is the norm in current RL literature. To address this, we introduce Frequency-Averaged Score (FAS) that is the normalized area under the curve (AUC) of the performance vs. decision frequency plot. We provide plots for all environments in the Appendix. We note that this experimental setup is similar to the challenge 7 introduced in by Dulac-Arnold et al. (2021) and SRL addresses the challenge of low throughput that is introduced in that work. The FAS captures the overall performance of the policy at different decision frequencies, timesteps or macro-action lengths. A High FAS indicates that the policy performance generalizes across:
  1. 1.
    Decision frequencies
  2. 2.
    Observation frequencies
  3. 3.
    Timestep sizes
@@
Figure 3: Avg. Reward vs FAS of different policies on stochastic timestep evaluations. Each marker is averaged over 10 episodes.
Figure 3 plots the performance of each policy vs. its individual FAS on the Ant environment.In all tested environments, except for the Inverted Double Pendulum, there is a strong Pearson correlation coefficient (greater than or equal to 0.82) between FAS and performance in stochastic conditions. This high correlation confirms the effectiveness of FAS as a metric for measuring a policy’s generalized performance across various timesteps and frequencies. The Inverted Double Pendulum, however, presents a unique challenge due to its requirement for high precision at low decision frequencies, leading to significantly lower FAS scores for all algorithms and thus it is an outlier. Comprehensive plots for all nine environments are included in the appendix (Fig. 10).
@@
We hypothesize that this limitation arises because GPM lacks mechanisms to address challenges associated with training on sequences of actions. For instance, altering the first action in a sequence can disrupt the optimality of subsequent actions, affecting the value function of deeper states and potentially causing deeper actions to diverge. SRL, on the other hand, mitigates this issue by incorporating a model and the ”temporal recall” mechanism, which help maintain consistency across the action sequence.
@@
Figure 5: Performance of SRL and model based online planning on different ASL. All markers are averaged over 5 trials. The error bars show standard error.
Table 4 compares the FAS score SRL to online planning using the same model in online planning versus the action sequences generated by the SRL policy. We see that SRL can learn action sequences and is competitive to model-based online planning. Notably, SRL performs better in environments with larger action and state space dimensions (See Figure 5 for an example of the Humanoid environment). Such environments are harder to model. Thus, SRL can leverage inaccurate models to learn accurate action sequences, further reducing the required computational complexity during training. We hypothesize that this superior performance is due to the fact that the actor learns a $J$-step action sequence concurrently, while online planning only produces one action at a time. Consequently, SRL is able to learn and produce long, coherent action sequences, whereas single-step predictions tend to drift, similar to the ’hallucination’ phenomenon observed in transformer-based language models.
6 Neural Basis for Sequence Learning
Unlike artificial RL agents, learning in the brain does not stop once an optimal solution has been found. During initial task learning, brain activity increases as expected, reflecting neural recruitment. However, after training and repetition, activity decreases as the brain develops more efficient representations of the action sequence, commonly referred to as muscle memory (Wiestler and Diedrichsen, 2013). This phenomenon is further supported by findings that sequence-specific activity in motor regions evolves based on the amount of training, demonstrating skill-specific efficiency and specialization over time (Wymbs and Grafton, 2015).
  1. 1.
    Basal ganglia (BG): Action chunking is a cognitive process by which individual actions are grouped into larger, more manageable units or ”chunks,” facilitating more efficient storage, retrieval, and execution with reduced cognitive load (Favila et al., 2024). Importantly, this mechanism allows the brain to perform extremely fast and precise sequences of actions that would be impossible if produced individually. The BG plays a crucial role in chunking, encoding entire behavioral action sequences as a single action (Jin et al., 2014; Favila et al., 2024; Jin and Costa, 2015; Berns and Sejnowski, 1996, 1998; Garr, 2019). Dysfunction in the BG is associated with deficits in action sequences and chunking in both animals (Doupe et al., 2005; Jin and Costa, 2010; Matamales et al., 2017) and humans (Phillips et al., 1995; Boyd et al., 2009; Favila et al., 2024). However, the neural basis for the compression of individual actions into sequences remains poorly understood.
  2. 2.
    Prefrontal cortex (PFC): The PFC is critical for the active unbinding and dismantling of action sequences to ensure behavioral flexibility and adaptability (Geissler et al., 2021). This suggests that action sequences are not merely learned through repetition; the PFC modifies these sequences based on context and task requirements. Recent research indicates that the PFC supports memory elaboration (Immink et al., 2021) and maintains temporal context information (Shahnazian et al., 2022) in action sequences. The prefrontal cortex receives inputs from the hippocampus.
  3. 3.
    Hippocampus (HC) replays neuronal activations of tasks during subsequent sleep at speeds six to seven times faster. This memory replay may explain the compression of slow actions into fast chunks. The replayed trajectories from the HC are consolidated into long-term cortical memories (Zielinski et al., 2020; Malerba et al., 2018). This phenomenon extends to the motor cortex, which replays motor patterns at accelerated speeds during sleep (Rubin et al., 2022).
7 Discussion Future Work
@@
The current RL framework encourages synchrony between the environment and the components of the agent. However, the brain utilizes components that act at different frequencies and yet is capable of robust and accurate control. SRL provides an approach to reconcile this difference between neuroscience and RL, while remaining competitive on current RL benchmarks. SRL offers substantial benefits over traditional RL algorithms, particularly in the context of autonomous agents in constrained settings. By enabling operation at slower observational frequencies and providing a gradual decay in performance with reduced input frequency, SRL addresses critical issues related to sensor failure and occlusion, and energy consumption. Additionally, SRL generates long sequences of actions from a single state, which can enhance the explainability of the policy and provide opportunities to override the policy early in case of safety concerns. SRL also learns a latent representation of the action sequence, which could be used in the future to interface with large language models for multi-modal explainability and even hierarchical reinforcement learning and transfer learning.
Future Work
Future work will incorporate bio-inspired features like attention mechanisms and knowledge transfer. Additionally, SRL can benefit from existing Model-Based RL approaches as it naturally learns a model of the world. In the Appendix, we demonstrate preliminary results of generative replay in the latent space. We believe that this is a promising direction to significantly improve upon the results in the paper.
In noiseless deterministic environments, a capable agent should achieve near-infinite horizon control for tasks like walking and hopping from a single state with minimal error corrections. Current approaches rely on external information at every state, which increases energy consumption and vulnerability to adversarial or missing inputs. Truly autonomous agents will need to implement multiple policies simultaneously, and simple tasks like walking can be performed with minimal input states if learned properly.
@@ A.2 Implementation Details
A.3 Practical Considerations on Low-Compute Hardware
@@ A.4 Learning Curves
A.5 Plots for Frequency Averaged Scores
@@
A.6 Plots for FAS vs. Stochastic Timestep Performance
@@
A.7 Generative Replay in Latent Space
@@
A.8 Clarification Figure
@@ A.9 Learning Curves by J
A.10 Randomized frame-skipping

### v4

@@ Abstract
Reinforcement learning (RL) is rapidly reaching and surpassing human-level control capabilities. However, state-of-the-art RL algorithms often require timesteps and reaction times significantly faster than human capabilities, which is impractical in real-world settings and typically necessitates specialized hardware. We introduce Sequence Reinforcement Learning (SRL), an RL algorithm designed to produce a sequence of actions for a given input state, enabling effective control at lower decision frequencies. SRL addresses the challenges of learning action sequences by employing both a model and an actor-critic architecture operating at different temporal scales. We propose a ”temporal recall” mechanism, where the critic uses the model to estimate intermediate states between primitive actions, providing a learning signal for each individual action within the sequence. Once training is complete, the actor can generate action sequences independently of the model, achieving model-free control at a slower frequency. We evaluate SRL on a suite of continuous control tasks, demonstrating that it achieves performance comparable to state-of-the-art algorithms while significantly reducing actor sample complexity. To better assess performance across varying decision frequencies, we introduce the Frequency-Averaged Score (FAS) metric. Our results show that SRL significantly outperforms traditional RL algorithms in terms of FAS, making it particularly suitable for applications requiring variable decision frequencies. Furthermore, we compare SRL with model-based online planning, showing that SRL achieves comparable FAS while leveraging the same model during training that online planners use for planning.
@@ 1 Introduction
Biological and artificial agents must learn behaviors that maximize rewards to thrive in complex environments. Reinforcement learning (RL), a class of algorithms inspired by animal behavior, facilitates this learning process (Sutton and Barto, 2018). The connection between neuroscience and RL is profound. The Temporal Difference (TD) error, a key concept in RL, effectively models the firing patterns of dopamine neurons in the midbrain (Schultz et al., 1997; Schultz, 2015; Cohen et al., 2012). Additionally, a longstanding goal of RL algorithms is to match and surpass human performance in control tasks (OpenAI et al., 2019; Schrittwieser et al., 2020; Kaufmann et al., 2023a; Wurman et al., 2022a; Vinyals et al., 2019; Mnih et al., 2015).
@@ 4 Sequence Reinforcement Learning
We introduce a novel reinforcement learning model capable of learning sequences of actions (macro-actions) by replaying memories at a finer temporal resolution than the action generation, utilizing a model of the environment during training. We provide the neural basis for our algorithm in the Appendix (A.9)
@@
5 Experiments
Overview
We evaluate our SRL approach on 11 continuous control tasks, comparing it against SAC (Haarnoja et al., 2019) and GPM (Zhang et al., 2022). We utilize the OpenAI Gym (Brockman et al., 2016) implementation of the MuJoCo environments (Todorov et al., 2012).
Experimental Setup
@@ 5 Experiments
Transitioning from simulation to real-world implementation (Sim2Real) in control systems is challenging because deployment introduces computational stochasticity, leading to variable sensor sampling rates (throughput) and inconsistent end-to-end delays from sensing to actuation (Sandha et al., 2021). This gap is not captured by the mean reward or return that is the norm in current RL literature. To address this, we introduce Frequency-Averaged Score (FAS) that is the normalized area under the curve (AUC) of the performance vs. decision frequency plot. We provide plots for all environments in the Appendix. We note that this experimental setup is similar to the challenge 7 introduced in by Dulac-Arnold et al. (2021) and SRL addresses the challenge of low throughput that is introduced in that work. The FAS captures the overall performance of the policy at different decision frequencies, timesteps or macro-action lengths. A High FAS indicates that the policy performance generalizes across decision frequencies, observation frequencies and timestep sizes.
@@
In all tested environments, except for the Inverted Double Pendulum, there is a strong Pearson correlation coefficient (greater than or equal to 0.82) between FAS and performance in stochastic conditions. This high correlation confirms the effectiveness of FAS as a metric for measuring a policy’s generalized performance across various timesteps and frequencies. The Inverted Double Pendulum, however, presents a unique challenge due to its requirement for high precision at low decision frequencies, leading to significantly lower FAS scores for all algorithms and thus it is an outlier. Comprehensive plots for all nine environments are included in the appendix (Fig. 7).
@@
Table 4 compares the FAS score SRL to online planning using the same model in online planning versus the action sequences generated by the SRL policy. We see that SRL can learn action sequences and is competitive to model-based online planning. Notably, SRL performs better in environments with larger action and state space dimensions. Such environments are harder to model. Thus, SRL can leverage inaccurate models to learn accurate action sequences, further reducing the required computational complexity during training. We hypothesize that this superior performance is due to the fact that the actor learns a $J$-step action sequence concurrently, while online planning only produces one action at a time. Consequently, SRL is able to learn and produce long, coherent action sequences, whereas single-step predictions tend to drift, similar to the ’hallucination’ phenomenon observed in transformer-based language models.
6 Discussion and Future Work
@@
The current RL framework encourages synchrony between the environment and the components of the agent. However, the brain utilizes components that act at different frequencies and yet is capable of robust and accurate control. SRL provides an approach to reconcile this difference between neuroscience and RL, while remaining competitive on current RL benchmarks. SRL offers substantial benefits over traditional RL algorithms, particularly in the context of autonomous agents in constrained settings. By enabling operation at slower observational frequencies and providing a gradual decay in performance with reduced input frequency, SRL addresses critical issues related to sensor failure and occlusion, and energy consumption. Additionally, SRL generates long sequences of actions from a single state, which can enhance the explainability of the policy and provide opportunities to override the policy early in case of safety concerns. SRL also learns a latent representation of the action sequence, which could be used in the future to interface with large language models for multimodal explainability and even hierarchical reinforcement learning and transfer learning.
@@ Appendix A Appendix
A. 1
SRL Algorithm A.1
A. 2
Hyperparameters A.2
A. 3
Implementation Details A.3
A. 4
Practical Considerations on Low-Compute Hardware A.4
A. 5
Learning Curves A.5
A. 6
Plots for Frequency Averaged Scores A.6
A. 7
Plots for FAS vs. Performance for Stochastic Timestep A.7
A. 8
Generative Replay in Latent Space A.8
A. 9
Neural Basis for Sequence Learning A.9
A. 10
Clarification Figure A.10
A. 11
Learning Curves by J A.11
A. 12
Randomized frame-skipping A.12
A. 13
Results for TempoRL Algorithm A.13
@@ A.3 Implementation Details
A.4 Practical Considerations on Low-Compute Hardware
@@ A.5 Learning Curves
A.6 Plots for Frequency Averaged Scores
@@
A.7 Plots for FAS vs. Performance for Stochastic Timestep
@@
A.8 Generative Replay in Latent Space
@@
A.9 Neural Basis for Sequence Learning
Unlike artificial RL agents, learning in the brain does not stop once an optimal solution has been found. During initial task learning, brain activity increases as expected, reflecting neural recruitment. However, after training and repetition, activity decreases as the brain develops more efficient representations of the action sequence, commonly referred to as muscle memory [Wiestler and Diedrichsen, 2013]. This phenomenon is further supported by findings that sequence-specific activity in motor regions evolves based on the amount of training, demonstrating skill-specific efficiency and specialization over time [Wymbs and Grafton, 2015].
  1. 1.
    Basal ganglia (BG): Action chunking is a cognitive process by which individual actions are grouped into larger, more manageable units or ”chunks,” facilitating more efficient storage, retrieval, and execution with reduced cognitive load [Favila et al., 2024]. Importantly, this mechanism allows the brain to perform extremely fast and precise sequences of actions that would be impossible if produced individually. The BG plays a crucial role in chunking, encoding entire behavioral action sequences as a single action [Jin et al., 2014, Favila et al., 2024, Jin and Costa, 2015, Berns and Sejnowski, 1996, 1998, Garr, 2019]. Dysfunction in the BG is associated with deficits in action sequences and chunking in both animals [Doupe et al., 2005, Jin and Costa, 2010, Matamales et al., 2017] and humans [Phillips et al., 1995, Boyd et al., 2009, Favila et al., 2024]. However, the neural basis for the compression of individual actions into sequences remains poorly understood.
  2. 2.
    Prefrontal cortex (PFC): The PFC is critical for the active unbinding and dismantling of action sequences to ensure behavioral flexibility and adaptability [Geissler et al., 2021]. This suggests that action sequences are not merely learned through repetition; the PFC modifies these sequences based on context and task requirements. Recent research indicates that the PFC supports memory elaboration [Immink et al., 2021] and maintains temporal context information [Shahnazian et al., 2022] in action sequences. The prefrontal cortex receives inputs from the hippocampus.
  3. 3.
    Hippocampus (HC) replays neuronal activations of tasks during subsequent sleep at speeds six to seven times faster. This memory replay may explain the compression of slow actions into fast chunks. The replayed trajectories from the HC are consolidated into long-term cortical memories [Zielinski et al., 2020, Malerba et al., 2018]. This phenomenon extends to the motor cortex, which replays motor patterns at accelerated speeds during sleep [Rubin et al., 2022].
A.10 Clarification Figure
@@ A.11 Learning Curves by J
A.12 Randomized frame-skipping